# Modeling, Analysis, and Optimization of Cascaded Power Amplifiers

### v2

@@ II System Model
We assume that the complex baseband equivalent of the input signal to the first PA in the cascade is denoted as $x_{n}^{(0)}$, with a power of $p_{0}$. PA behavior under normal operation conditions, in the simplest case, can be modeled using the third-order memoryless polynomial model [3]. Thus, the output signal $y_{n}^{(k)}$ of the $k$th PA can be described as
@@ III-C Analysis of the Approximating Expressions
From equations (7)–(11), it can be observed that the amount of distortions at the output of the cascade of $K$ PAs strongly depend on their parameters $g^{(k)}$ and $\alpha^{(k)}$, $k=1,\dots,K$, and the power $p_{0}$ of the input signal $x_{n}^{(0)}$ to the first stage. Specifically, a linear combination of $\alpha^{(k)}$ in (9) relies on the product of quadratic terms $(g^{(k)})^{2}$, and noise level in (11) is influenced by the product of the linear terms $g^{(k)}$. The coefficient $\alpha^{(k)}$ represents the inherent nonlinear distortion characteristics of the $k$th amplifier, that is generally difficult to change without redesigning the amplifier. In contrast, the input power $p_{0}$ and linear gain $g^{(k)}$ can be more easily adjusted in practice. In this paper, we consider that the PAs in cascade have the same nonlinear characteristics, i.e., $\alpha^{(k)}=\alpha$. Thus, varying the parameters $p_{0}$ and $g^{(k)}$ can lead to minimizing the total nonlinearities at the output of the $K$th PA.
@@ V-A The Choice of the Starting Point for the Optimization
Figure 2: AM/AM of two initial scenarios of the cascaded PAs for $\text{SNR}=50$ dB.
Figure 3: PSD of two initial scenarios of the cascaded PAs for $\text{SNR}=50$ dB.
@@
Input power opt. (Scenario 2) 0.53 0.27 0.14 0.07 0.03
Input power and equal PA gains optimization 0.64 0.40 0.25 0.24 0.24
Input power and unequal PA gains optimization 0.64 0.40 0.33 0.37 0.38
@@
Equal PA gains optimization $1$ 1.06 1.06 1.06 1.05 1.05
$2$ 1.06 1.06 1.05 1.05
$3$ 1.06 1.05 1.05
$4$ 1.05 1.05
$5$ 1.05
Unequal PA gains optimization $1$ 1.06 0.84 0.70 0.70 0.70
$2$ 1.30 1.23 0.97 0.90
$3$ 1.30 1.30 1.09
@@
Input power and equal PA gains optimization $1$ 1.30 1.30 1.29 1.23 1.18
$2$ 1.30 1.29 1.23 1.18
$3$ 1.29 1.23 1.18
$4$ 1.23 1.18
$5$ 1.18
Input power and unequal PA gains optimization $1$ 1.30 1.30 1.11 1.02 1.03
$2$ 1.30 1.30 1.06 1.03
$3$ 1.30 1.30 1.03
@@ V-B3 Joint Input Power and PA Gains Optimization
Finally, joint optimization of the PA gains and input power is considered here. Scenario 1 as the starting point and $\varepsilon=0.3$ are also chosen here. The results of the optimization are shown in Figs. 46, and the optimized values of the input signal level and PA gains are listed in Tables II and II, respectively. One can observe that incorporating the input power level into the optimization, along with the gains, further enhances the potential for reducing overall distortions and mitigating the severe effects of cascading $K$ PAs, however, the difference between equal and unequal gains is small and becomes noticeable only for $K\geq 4$.
Figure 4: Results of the optimization of the cascaded PAs parameters in terms of NMSE and ACLR for $\text{SNR}=50$ dB.
Figure 5: AM/AM of the output after $K$ PAs with the optimized input power and PA gains (equal and unequal gains), $\text{SNR}=50$ dB.
Figure 6: PSD of the output after $K$ PAs with the optimized input power and PA gains (equal and unequal gains), $\text{SNR}=50$ dB.
@@ VI Conclusions
This paper introduced a model of cascaded PAs and provided an analysis of the total nonlinearities, which accumulate as the number of PAs increases. The NLS algorithm was proposed to optimize PA gains and input power levels, thereby minimizing the overall impact of cascaded nonlinearities. Simulation results demonstrated the performance improvement of cascaded structures optimized using the proposed framework.

### v3

@@ II System Model
We assume that the complex baseband equivalent of the input signal to the first PA in the cascade is denoted by $x_{n}^{(0)}$, with power of $p_{0}$. PA behavior under normal operation conditions, in the simplest case, can be modeled using the third-order memoryless polynomial model [3]. Thus, the output signal $y_{n}^{(k)}$ of the $k$th PA can be described as
@@ III-C Analysis of the Approximating Expressions
From equations (7)–(11), it can be observed that the amount of distortions at the output of the cascade of $K$ PAs strongly depend on their parameters $g^{(k)}$ and $\alpha^{(k)}$, $k=1,\dots,K$, and the values of the input signal $x_{n}^{(0)}$ to the first stage, i.e., the power $p_{0}$. Specifically, a linear combination of $\alpha^{(k)}$ in (9) relies on the product of quadratic terms $(g^{(k)})^{2}$, and noise level in (11) is influenced by the product of the linear terms $g^{(k)}$. The coefficient $\alpha^{(k)}$ represents the inherent nonlinear distortion characteristics of the $k$th amplifier, that is generally difficult to change without redesigning the amplifier. In contrast, the input power $p_{0}$ and linear gain $g^{(k)}$ can be more easily adjusted in practice. In this paper, we consider that the PAs in cascade have the same nonlinear characteristics, i.e., $\alpha^{(k)}=\alpha$. Thus, varying the parameters $p_{0}$ and $g^{(k)}$ can lead to minimizing the total nonlinearities at the output of the $K$th PA.
@@ V-A The Choice of the Starting Point for the Optimization
Figure 2: AM/AM of the cascaded PAs for the two initial scenarios.
Figure 3: PSD of the cascaded PAs for the two initial scenarios.
@@
Input power opt. (Scenario 2) 0.49 0.23 0.11 0.05 0.02
Input power and equal PA gains optimization 0.65 0.40 0.25 0.19 0.19
Input power and unequal PA gains optimization 0.65 0.40 0.27 0.30 0.32
@@
Equal PA gains optimization $1$ 1.07 1.07 1.07 1.06 1.06
$2$ 1.07 1.07 1.06 1.06
$3$ 1.07 1.06 1.06
$4$ 1.06 1.06
$5$ 1.06
Unequal PA gains optimization $1$ 1.07 0.86 0.70 0.70 0.70
$2$ 1.30 1.26 0.99 0.82
$3$ 1.30 1.30 1.21
@@
Input power and equal PA gains optimization $1$ 1.30 1.30 1.30 1.26 1.21
$2$ 1.30 1.30 1.26 1.21
$3$ 1.30 1.26 1.21
$4$ 1.26 1.21
$5$ 1.21
Input power and unequal PA gains optimization $1$ 1.30 1.30 1.24 1.02 1.02
$2$ 1.30 1.30 1.17 1.03
$3$ 1.30 1.30 1.13
@@ V-B3 Joint Input Power and PA Gains Optimization
Finally, joint optimization of the PA gains and input power is considered here. Scenario 1 as the starting point and $\varepsilon=0.3$ are also chosen here. The results of the optimization are shown in Figs. 66, and the optimized values of the input signal power and PA gains are listed in Tables II and II, respectively. One can observe that incorporating the input power level into the optimization, along with the gains, further enhances the potential for reducing overall distortions and mitigating the severe effects of cascading $K$ PAs, however, the difference between equal and unequal gains is small and becomes noticeable only for $K\geq 4$.
Figure 4: Results of the optimization of the cascaded PAs parameters in terms of NMSE and ACLR.
Figure 5: AM/AM of the output after $K$ PAs with the optimized input power and PA gains (equal and unequal gains).
Figure 6: PSD of the output after $K$ PAs with the optimized input power and PA gains (equal and unequal gains).
@@ VI Conclusions
This paper studied modeling of cascaded PAs and provided an analysis of the total nonlinearities, which accumulate as the number of PAs increases. The NLS algorithm was proposed to optimize PA gains and input power levels, thereby minimizing the overall impact of cascaded nonlinearities. Simulation results demonstrated the performance improvement of cascaded structures optimized using the proposed framework.